OpenStudy (anonymous):

5-i, 6,7-i... squrt5,sqrt10,2sqrt5... 1,1,2,3,5,8 find the indicated s sub n

OpenStudy (campbell_st):

the 1st sequence seems to be \[5 - i^n\] where n is the term number the 2nd is \[\sqrt{5},\sqrt{10},\sqrt{20}...\] this looks like \[\sqrt{n \times5}\] where n is the term number last one is a fibonacci sequence.... easy to google the method...

OpenStudy (campbell_st):

oops the square root sequence is \[\sqrt{5} \times \sqrt{2^{n-1}}\] where n is the term number